Problem 5 is related to Lecture 16 Problem 5. SPT and WSPT (1.3 points) (a) The table below displays the processing time for 6 projects Project P1 P2 P3 P4 P5 P6 Processing time (hours) 2.5 4 Please use the Shortest Processing Time (SPT) rule to sequence the 8 projects. (0.2 points) P2 P5 ProjectP1 Sequenc P3 P4 P6 Using the Shortest Processing Time rule, the average flow time of these 6 project hours Please show detailed analysis below (0.3 points): (b) Processing times and weights for five jobs are shown in the table below P5 Job Processing Time (minutes) Weights P1 P2 P3 P4 2.5 Please use the Weighted Shortest Processing Time (WSPT) rule to sequence the 5 jobs. (0.3 points) P1 P2 P3 P4 P5 Sequenc Show detailed analysis for your sequence below:
(c) Processing times are shown in the table below. The weight for each job is intangible. The importance relations among these 4 jobs are as below: P1 and P2 are equally important. P3 is twice as important as P1 P4 is three time important as P1 Please use these relations to assign weight to each job and use WSPT to sequence the 4 jobs. (0.5 points) Job P1 P2 P3 P4 Processing Time (minutes) 4 2.5 Weight Job P1 P2 P3 P4 Sequenc Show your analysis for the sequence below:
Answer #1

a.

According to SPT rule the following sequence is applicable:

(schedule job with shortest processing time first, sequence the jobs in increasing order of processing time)

1st

2nd

3rd

4th

5th

6th

P1

P4

P3

P2

P5

P6

First determine Job start time = 0 (for first job) or Fi-1 for the jobs second onwards

Determine finish start time = Start time of job (Si) + Processing time (Pi) of job

Job Sequence

Processing time

Job start time

Job Finish Time

Job

pi

Si

Fi = Si + Pi

P1

1

0

1

P4

2

1

1+2

= 3

P3

2.5

3

3+2.5

= 5.5

P2

3

6

8.5

P5

4

9

12.5

P6

5

13

17.5

Total

17.5

48

Flow Time

Total flow time = 48

No. of Jobs = 8

Average Flow time = Total Flow time/no. of jobs = 48 / 6 = 8 hours

b.

Weighted Shortest Processing time Rule

1)determine ration of weight (Wj) to processing time (Pj)

2)arrange the jobs in decreasing order of Wj / Pj ratio (highest ratio job should be scheduled first)

Jobs (j)

Processing Time (Pj)

Weight (Wj)

(Wj) / (Pj)

Sequence

P1

1

2

2

1st

P2

3

1

0.333333

5th

P3

2.5

1.5

0.6

3rd

P4

2

3

1.5

2nd

P5

5

2

0.4

4th

Job Sequence

Processing time

Job start time

Job Finish Time

Job

pi

Si

Fi = Si + Pi

P1

1

0

1.000

P4

2

1

3.000

P3

2.5

3

5.500

P5

5

6

10.500

P2

3

11

13.500

Total

13.5

33.500

Total flow time = 33.5

No. of jobs = 5

Average flow time = total flow time/No. of jobs = 33.5/5 = 6.70